OpenStudy (anonymous):

Simplify (4a -1) 2+a

Parth (parthkohli):

Is that \((4 a - 1)^2 + a\)?

OpenStudy (anonymous):

no exponents

Parth (parthkohli):

\[(4 a - 1)(2 + a) \]?

OpenStudy (anonymous):

yea but ( 2 +a has no brackets

Parth (parthkohli):

Then I don't think that the question makes sense. Does your question mean \(2(4a - 1) + a\)?

OpenStudy (anonymous):

You have another distributive property problem, but then just add "a" to the result. You can first do 2(4a-1), then add "a" to that result. Does that make sense?

OpenStudy (anonymous):

@rubie the answer is 16a^2-7a+1

OpenStudy (anonymous):

@Math4Life, the problem is multiply by 2, nothing is squared.

OpenStudy (anonymous):

(4a-1)2 = 2(4a-1) Use the distributive property to multiply, then add "a" to the result

OpenStudy (anonymous):

@catherine oh i didnt see that. Then 4a^2+7a-2

OpenStudy (anonymous):

@Math4Life, giving someone an answer is not helpful, especially if it's a wrong answer

OpenStudy (anonymous):

@catherine thats not the wrong answer..

OpenStudy (anonymous):

It sais to simplify.

OpenStudy (anonymous):

Simplified its 4a^2+7a-2

OpenStudy (anonymous):

No it's not, nothing is squared.

OpenStudy (anonymous):

@catherine you know nothing about math.. i'm in AP calculus and i'm a freshman...

OpenStudy (anonymous):

@catherine it's not even hard to understand how they come together and make ^2

OpenStudy (anonymous):

(4a -1) 2+a Solution with steps: (4a -1) 2+a = 2(4a -1)+a (Commutative Prop.) = (2*4a - 2*1) + a (Distributive Prop.) = (8a - 2) + a (Multiply) = 9a - 2 (Simplify by combining like terms)

OpenStudy (anonymous):

@Math4Life Do you see your mistake now?

OpenStudy (anonymous):

@catherine ah i'm sorry for acting like a jerk lol

OpenStudy (anonymous):

@Math4Life And, yes, I do know math. My degree was Mathematics, and I was a high school math teacher. But my point is, don't think you are helping someone by giving out answers. At least when I gave the answer, I also gave every step with a reason for that step. The purpose of study groups is to give help, not answers.